OpenStudy (anonymous):

Train traveling on parrellel tracks in same direction. Train A at 60mph passes station at 11.15. Train B at 70mph passes same station at 11:27. At what time will train b catch up with train A.

OpenStudy (anonymous):

60x+12=70x; 12=10x; x=12/10=1.2; train b will catch up in 1.2 mins

OpenStudy (anonymous):

that's not right

OpenStudy (anonymous):

60x+720=70x 720=10x x=72 minutes

OpenStudy (anonymous):

I got 72 minutes but it seems a very long time when it is only 12 minutes behind train a.

OpenStudy (anonymous):

Nvm, I found my error

OpenStudy (anonymous):

Train A= 60x+(12/60)*60 (it is because Train A is 12 minutes ahead --> 12 minutes is 12/60 hour) Train B= 70x 60x+12=70x 12=10x x=1.2 minutes

OpenStudy (anonymous):

you were right, gj

OpenStudy (anonymous):

so it is not 60(x+12)=70x its just 60x+12=70x;12=10x;x=12/10;1.2 minutes

OpenStudy (anonymous):

yes, since we are working with unit hours, we have to use hour--> 12 minutes=12/60 hours

OpenStudy (anonymous):

so will i add 12 minutes or 1.2 minutes

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

since the train B will pass Train A in 1.2 minutes, 11:27 + 1.2

OpenStudy (anonymous):

the question is when will train b catch up with train a. will you add 12minutes or 1.2 minutes to 11:27

OpenStudy (anonymous):

ok thanks and have a great day.